OpenStudy (anonymous):

Find the area of an isosceles triangle with a base of length 12 cm and congruent of length 10 cm.

OpenStudy (anonymous):

use Herons' Formula... Do you know this formula??????

OpenStudy (anonymous):

12^2+b^2=10^2

OpenStudy (anonymous):

Why r u using Pythagoras theorem here????? it is valid only for right angled triangles...

OpenStudy (anonymous):

according to Heron's formula area of a triangle = sqroot[s(s-a)(s-b)(s-c) a+b+c where s = ------- 2 and a, b and c are the lengths of the three sides of the triangle

OpenStudy (anonymous):

From the figure u hv sent, it appears we first hv to find the height using the Pythagoras theorem In an isosceles triangle, the perpendicular from the vertex bisects the opposite side into two equal pcs of 6 cm each So in the Right angles triangle we have Hypotenuse = 10cm and base = 6cm (Hypotenuse)^2 = (Base)^2 + (Perpendicular)^2 Here it becomes 10^2 = 6^2 + (P)^2 100 = 36 + (P)^2 100-36 = (P)^2 64=(P)^2 root64 = P 8 =P So u hv height 8 cm Area = 1/2 x base x height = 1/2 x 12 cm x 8 cm area = 96/2 cm^2 = 48 cm^2